If the effective mass of electron in copper is m∗=1.01 me(me = mass of an electron) and the electric conductivity of copper is 5.76 ×107 Ω−1 m−1 with EF=7 eV then calculate the Fermi mean free path for copper.
1
Expert's answer
2021-09-20T10:01:05-0400

Find the relaxation time:


"\tau=\frac{m^*\gamma}{q^2n}."

Fermi velocity:


"v_F=\sqrt{\frac{2E_F}{m^*}}."


The Fermi mean free path for copper is


"\Lambda=\nu_F\tau.\\\space\\
\Lambda=\sqrt{\frac{2E_F}{m^*}}·\frac{m^*\gamma}{q^2n}=\frac{\gamma\sqrt{2E_Fm^*}}{q^2n},\\\space\\
\Lambda=3.22·10^{21}\text{ m}."